# Uniform Field in Microwave Cavities Through the Use of Effective Magnetic Walls

### v1

@@ Abstract
Our experimental results show improved field homogeneity due to the optimized gap between the WM sample and the resonator walls. The slight discrepancies between experimental and simulation results, seen in the field maps of Fig. 8, are likely due to experimental limitations. Variations in wire length, caused by manual soldering, and potential material deformations from excessive heat may have led to wire or PCB plate bending. Additionally, the scanning procedure, where the antenna moved through holes, could have introduced slight variations if the antenna was not perfectly centered at each hole. This may explain the difficulty in pinpointing the exact location of the maximum field, which did not align precisely with the central hole as simulations predicted.
In conclusion, we introduced a method to enhance field homogeneity in WM resonators by increasing the air gap between the resonator walls and the WM sample, with the field homogeneity peaking at a gap of one-quarter of the plasma wavelength. Our theoretical model and simulations predict a form factor of 0.92 for the effective WM model and 0.89 for the physical wire array, with the resonant mode frequency shifting toward the plasma frequency. Experimental validation was carried out with two prototype resonators: a regular WM resonator and an optimized one with a $\lambda_{p}/4.2$ shift. We measured the field distribution and form factor, finding significant improvements in the form factor, from 0.61 to 0.77 experimentally, and from 0.64 to 0.80 in simulations. These results validate the use of the TM000 mode for improving the form factor in plasma haloscopes. Future work will focus on advancing cavity design, including exploring diverse cross-sections and tunable resonance frequencies.
@@ Acknowledgements
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.

### v3

@@ III Experiment
So, our experimental results demonstrate an improvement in field homogeneity due to the optimized gap between the WM sample and the resonator walls. Slight discrepancies between experimental and simulation results seen in the field maps of Fig. 8 can be attributed to experimental limitations in the prototype fabrication and the field scanning procedure. Due to manual soldering of the wires, variations in wire length within the prototypes may exist. Additionally, excessive heat during soldering could have caused material deformations, leading to bending in either the wires or PCB plates. Moreover, the scanning procedure, involving the antenna traversing through holes, might have introduced slight variations in field measurements if the antenna was not perfectly centered at each hole. This could explain the difficulty in differentiating between high field levels and the exact location of the maximum field, which was not precisely at the central hole as simulations predicted.
IV Conclusions
We introduced a new method to enhance field homogeneity within WM resonators. Our theoretical model, supported by numerical simulations, demonstrated that properly increasing the air gap between the resonator walls and the WM sample, one can effectively manipulate the field homogeneity. When the air gap reached one-quarter of the resonant wavelength, the field homogeneity peaked, theoretically resulting in a form factor of 0.92 for an effective WM model (theory and simulations) and 0.89 for a physical array of wires (simulations). Additionally, this enhancement in the form factor was accompanied by a shift in the resonant mode frequency toward the plasma frequency.
We experimentally validated the proposed method by developing two prototype resonators: a regular WM resonator containing the WM sample with the conventional half-period air gap between the wires and the walls and an optimized $\lambda_{p}/4.2$-shifted resonator. We devised a method to measure the field distribution within the transverse section of the cavities, enabling the subsequent calculation of the form factor. This method involved measuring the magnitude of the $S_{21}$ parameter using a fixed antenna at one end of the resonator and a scanning antenna at the other end. The scanning antenna traversed predrilled holes within the resonator. Comparing experimental results with full-wave simulations, we found that the form factor in a standard WM resonator was 0.61 (experiment) and 0.64 (simulation), whereas in the approximately $\lambda/4.2$-shifted WM resonator, the form factor increased to 0.77 (experiment) and 0.80 (simulation). Our findings validate and substantiate using the TM000 mode to improve the form factor in plasma haloscopes, opening new avenues for WM cavity applications where precise control of the internal field profile is essential.
In order to realize a practical experiment, there are two main avenues for future work. The first is to adapt the cavity design to have a cross-section better suited for a cylindrical bore of a solenoidal magnet. Secondly, given purpose of the plasma haloscopes inherently requires their resonance frequency to be tunable in order to search for dark matter at different frequencies, we plan to investigate the effectiveness of the proposed approach in improving the form factor when the mode frequency varies. This will involve considering both operation away from the quarter-wave condition and adjustment of the walls to maintain it, similarly to the tuneable rectangular cavity in Ref. [28].
@@ Acknowledgements
The authors thank Eugine Koreshin for helpful discussions and also thank the members of the ALPHA Consortium for discussion and support.